# mybibliography.bib

### v1

@@ 2.3 Choice of embeddings
where $c>0,\sigma_{2}>\sigma_{1}>0,$ will not degenerate, and allow effective training of the embedding.
This has the detriment of adding $n\cdot\operatorname{dim}\mathcal{M}$ trained parameters to the model. However, $\operatorname{dim}\mathcal{M}$ is in practice always substantially smaller than $n$, hence the additional trained data is a small fraction of the trained data of the RNN. Additional computations are also necessary for the recalculation of the regularizing terms, however: in our experiments, we found that training the embedding resulted in substantial increases to training time. In practice it is not necessary for the embeddings to be trained: the learned update vectors will naturally adapt to randomly chosen embeddings.
@@ 3 Limitations
One flaw of our approach is that the moduli space $\mathcal{M}$ is chosen as a hyperparameter of training, rather than adjusting to the input data. In principle it is possible for trained embeddings (Section 2.3) with $\mathcal{M}=\mathbb{R}^{n}$ or $\mathcal{M}=S^{n}$ to allow moduli spaces to be discovered simultaneously with training, but our preliminary explorations were of mixed success. Sample results from these experiments are shown in Table 3 and in Appendix H). Creating a modified training arc which allows the moduli space to be learned and detected is an interesting avenue for future research. One possible conflating factor for our experiments is our sparsification scheme—magnitude pruning. Magnitude pruning is the simplest pruning method available, and we chose it in order to simplify the discussion and avoid conflating factors in our experiments. Changes to this sparsification method might impact the level of improvement moduli regularization offers, as compared to alternative regularization schemes, though our lottery ticket hypothesis [frankle2018lottery] experiments (Section 4) will remain of interest in this comparison. Finally, due to the computational costs, we did not extensively optimize the hyperparameters in our experiments, and tested on limited architectures.
@@ 4 Results
We test the proposed moduli regularization (Equation (5)) on two benchmark tasks: navigation and natural language processing (NLP). We study both tasks using the circle, the sphere, the Klein bottle, the 2-torus, and the 6-torus as moduli spaces, with the parameterizations described in Appendix D, and with the difference of Gaussians inhibitor function (Equation (7)). Additional data comparing other inhibitor functions is shown in Appendix E.
We chose NLP as a benchmark because it is of great general interest and is one of the most common applications of RNNs. Our original interest, however, was in the structures underlying navigation in RNNs, and navigation simulations [burak2009accurate, gardner2022toroidal, guanella2007model, schaeffer2023self, sorscher2023unified] provided inspiration for our constructions. This set of benchmarks provides a useful contrast: navigation is a geometric regression task, with known moduli spaces on which the recurrent dynamics occur in both mammalian brains [burak2009accurate, gardner2022toroidal, guanella2007model] and simulations [schaeffer2023self, sorscher2023unified]. NLP is a classification task which lacks a clean geometric description [elman1991distributed]. Despite this distinction, we show that moduli regularization, particularly along the torus and the Klein bottle with difference of Gaussians inhibitors, is a highly effective technique for model sparsification for both benchmarks.
@@
The lottery ticket hypothesis [frankle2018lottery] suggests that sparse architectures arise from fortunate weight initializations. In other words, the training process for sparse neural architectures is unstable: the subspace of potential weight initializations which will cause training a model to converge to a high-quality net is low measure in the space of all potential weight initializations. This contrasts with dense neural nets, which are robust to random weight initializations. As a result, retraining models on a sparse architecture without using the original weight initialization deteriorates performance [frankle2018lottery, gale2019state]. We show that the sparse structures learned with moduli regularization are more stable than the sparse structures learned without it: models retrained on moduli-generated sparse architectures outperform models retrained on traditional sparse architectures. In several of our NLP experiments, we even achieve superior results after reinitialization and retraining ab initio sparse models as opposed to the original sparsified models.
@@ 4.1 Navigation
We study a single-layer RNN designed for autonomous landmark mapping and navigation, using the architecture of [sorscher2023unified]. The RNN initializes its hidden state with a two-dimensional position, then takes as input a sequence of two-dimensional velocity vectors. The RNN is trained to identify which landmarks are nearest to the final position. Training data in the form of trajectories can be freely generated for the RNN, which alleviates the problem of overfitting. This allows us to study the effect of regularization on the sparsity of the underlying model, isolated from the other benefits of regularization. Hyperparameters for our experiments are listed in Appendix A.1. Trajectories occur in a 2.2 m square box, in which untrained models have a mean error of $\approx 100$ cm.
@@
Baseline 4.79 (0.10) 4.67 (0.07) 5.85 (0.09) 4.77 (0.09)
75% Sparse 4.60 (0.06) 4.86 (0.30) 6.59 (0.85) 34.78 (42.10)
75% Lottery 8.20 (4.16) 7.28 (0.69) 35.44 (39.83) 60.37 (35.54)
90% Sparse 5.87 (0.55) 5.69 (0.43) 5.81 (0.45) 67.42 (33.27)
90% Lottery 15.72 (3.94) 12.49 (2.03) 20.67 (5.68) 46.56 (7.69)
Lottery tickets: We test the lottery ticket hypothesis [frankle2018lottery] with moduli regularizers (Tables 1 and 2) at 75% and 90% sparsity. The sparse networks that arise from moduli regularizers consistently outperform the alternatives with ab initio sparsity and random initializations: compare especially the columns ‘Torus’ and ‘Shuffled’ in Table 1.
Sensitivity analysis on the regularizing factor ($\lambda$ in Equation (4)) can be found in Appendix G. This RNN architecture was originally introduced by [sorscher2023unified] to study grid cells, certain specialized neurons in the entorhinal cortex which provide an internal map of space [giocomo2011computational, moser2008place, whittington2022build]. After the discovery of grid cells in 2005 [hafting2005microstructure], several authors [burak2009accurate, guanella2007model] show that toroidal neural connectivity can give rise to grid cell structures. However, to the best of our knowledge, no connectivity structures outside of the torus and $\mathbb{R}^{2}$ have been seriously investigated—the hand construction of these networks is time consuming, and small perturbations of weights can damage their effectiveness [vafidis2022learning]. One surprising application of our moduli regularization is then the great ease with which it shows the learning advantages of the torus over arbitrary spaces, such as the sphere, for neural attractor dynamics on navigation. The success of toroidal, circular, and Klein bottle sparsification suggests that moduli regularization may provide a useful lens for exploratory work on highly structured neural connectivity. We expected these regularizers to outperform the sphere and 6-torus because $\mathbb{R}^{2}$ is a covering space (see summary in Appendix C) for the Klein bottle and the torus, making them apt compact representations for two-dimensional space. Since the torus is isomorphic to the product of two circles, a circular moduli space could also plausibly recreate toroidal connectivity.
4.2 Natural language processing
Natural language processing lacks clear a priori geometry to exploit [elman1991distributed]. Indeed, local stability of the hidden state is not a common hypothesis for NLP architectures like long-short term memory [hochreiter1997long]. We test an RNN with three recurrent layers trained to predict subsequent words in a sequence on the Wikitext-2 dataset of high-quality Wikipedia articles [merity2016pointer]. Our implementation is based on the implementation of [pytorch_examples], which is a representative implementation of natural language RNNs. Hyperparameters for our experiments are listed in Appendix A.2.
Table 4: Mean test accuracy (standard deviation) of next-word prediction RNNs at baseline, 90%, and 98% sparsity, and testing the lottery ticket hypothesis with trained sparse architectures.
Torus No reg. $L_{1}$ Shuffled
Baseline 8.02% (1.73) 2.34% (0.16) 2.29% (0.03) 9.38% (0.49)
90% Sparse 10.31% (1.33) 5.30% (3.37) 2.85% (0.00) 8.44% (3.11)
90% Lottery 9.77% (0.81) 2.85% (0.00) 2.79% (0.11) 7.42% (2.49)
98% Sparse 9.59% (1.07) 2.85% (0.00) 3.78% (3.51) 9.41% (1.28)
98% Lottery 9.17% (0.51) 2.76% (0.17) 2.85% (0.00) 9.05% (0.49)
@@ 4.2 Natural language processing
Table 5: Mean test accuracy (standard deviation) of next-word prediction RNNs at baseline, 90%, and 98% sparsity, trained with a variety of moduli regularizers, and testing the lottery ticket hypothesis. Underlined values are ab initio sparse models that outperform their sparsified antecedents.
@@
Baseline 8.02% (1.73) 9.06 (1.56) 8.25 (0.95) 8.34 (1.03) 8.88 (1.26)
90% Sparse 10.31% (1.33) 9.82 (1.29) 9.41 (1.47) 9.71 (0.98) 9.83 (0.76)
90% Lottery 9.77% (0.81) 10.07 (1.15) 9.57 (1.12) 9.51 (0.98) 10.60 (0.64)
98% Sparse 9.59% (1.07) 10.19 (0.69) 9.36 (1.06) 8.91 (0.46) 9.26 (0.86)
98% Lottery 9.17% (0.51) 8.85 (0.34) 9.48 (0.56) 10.09 (0.21) 8.68 (0.56)
Lottery tickets: We test the lottery ticket hypothesis [frankle2018lottery] for each experiment. Our results conform with the hypothesis across control experiments, indicating that the learned sparse architectures are unstable. This is also true for some—but not all—tested moduli regularizers. Several models trained with moduli regularizers improved after randomizing the original initialization (Circle, Sphere, Klein bottle, 6-Torus in Table 5), suggesting these architectures are stable with respect to weight initialization. Lacking knowledge of the true moduli space of the system results in no serious improvements compared to random sparsification (the Shuffled column of Table 4). However, the broad stability of sparse architectures produced with moduli sparsification suggests that geometric systems are more stable than non-geometric ones.
@@
Our results show that moduli regularization remains a highly effective sparsification method in non-geometric RNNs. Moduli regularizers (Table 5) improve the accuracy of trained NLP models compared to both non-regularized and $L_{1}$ regularized models (Table 4). While they do not consistently outperform random sparsification at 98% sparsity, some moduli regularizers create sparse architectures that are more resilient to retraining with randomized weights, in some cases defying the lottery ticket hypothesis [frankle2018lottery].

### v2

@@ 2.3 Learning the moduli space
where $c>0,\sigma_{2}>\sigma_{1}>0,$ or the Ricker wavelet (Equation (7)) will not degenerate, and allow effective training of the embedding.
Learning the moduli regularizers comes at the detriment of adding $n\cdot\operatorname{dim}\mathcal{M}$ trained parameters to the model. However, $\operatorname{dim}\mathcal{M}$ is in practice substantially smaller than $n$, hence the additional trained data is a small fraction of the trained data of the RNN. Additional overhead computations are also necessary for the recalculation of the regularizing terms, requiring $O(n^{2})$ computations at the end of gradient step epoch. In practice, we found this overhead was negligible when $n<500$, and small when $n<1000$.
@@ 3 Limitations
Moduli regularization shows some sensitivity to the regularizing factor, a common limitation for regularization techniques. The choice of moduli space $\mathcal{M}$ and inhibitor $f$ are further additional hyperparameter which can be hard to properly select, and are slow to be learned when working with RNNs with very high dimensional hidden spaces. Interesting avenues for future research include choosing a moduli space using manifold discovery, a wide range of techniques for storing data and computing optimal, nonlinear representation schemes [7, 17, 24, 42, 48].
One possible conflating factor for our experiments is our sparsification scheme—magnitude pruning. Magnitude pruning is the simplest pruning method available, and we chose it in order to simplify the discussion and avoid conflating factors in our experiments. Changes to this sparsification method might impact the level of improvement moduli regularization offers, as compared to alternative regularization schemes, though our lottery ticket hypothesis [9] experiments (Section 4) will remain of interest in this comparison. Finally, due to the computational costs, we did not extensively optimize the hyperparameters in our experiments, and tested on limited architectures.
@@ 4 Results
We test moduli regularization (Equation (6)) on three benchmark tasks: navigation, natural language processing (NLP), and the adding problem. Navigation is a geometric regression task, with known moduli spaces on which the recurrent dynamics occur in both mammalian brains [2, 14, 18] and simulations [4, 43, 49]. NLP is a classification task which lacks clean geometric descriptions [6]. The adding problem is an experiment designed to study the ability for RNNs to maintain long term memories, and likewise lacks obvious geometry. Despite this distinction, we show that moduli regularization is a highly effective technique for model sparsification for all benchmarks.
We study these tasks using both trained and untrained moduli regularizers in a variety of low dimensional ambient spaces $\mathcal{M}$ for comparison. We find that training the regularizer, assuming lack of knowledge of a good moduli space, is highly effective for both navigation and adding problem experiments. In our NLP experiments, however, training the regularizer seemed to enable additional overfitting to occur, as opposed to fixing random embeddings.
@@
The lottery ticket hypothesis [9] suggests that sparse architectures arise from fortunate weight initializations. In other words, the training process for sparse neural architectures is unstable: the subspace of potential weight initializations which will cause training a model to converge to a high-quality net is low measure in the space of all potential weight initializations. This contrasts with dense neural nets, which are robust to random weight initializations. As a result, retraining models on a sparse architecture without using the original weight initialization deteriorates performance [9, 11]. We show that the sparse structures learned with moduli regularization are more stable than the sparse structures learned without it: models retrained on moduli-generated sparse architectures outperform models retrained on traditional sparse architectures. In several of our NLP experiments, we even achieve superior results after reinitialization and retraining ab initio sparse models as opposed to the original sparsified models. For all experiments, we report both the error of progressively sparsified models, labeled $X$Sparse, and the results of retraining with randomized weights in the learned sparse architecture, labeled $X$Lottery. Lottery ticket experiments were additionally conducted with the same regularizer as was used in the antecedent model. When using trained embeddings, we ceased training embeddings for lottery ticket experiments (Tables 1–9).
@@ 4.1 Navigation
We study a single-layer RNN designed for autonomous landmark mapping and navigation, using the architecture of [4]. A 2.2 m square arena is fixed, and $\ell$ uniformly randomly chosen points in the arena are fixed throughout training, and referred to as landmarks. The RNN initializes its hidden state with a two-dimensional position $x_{0}$, which is transformed to $\operatorname{softmax}(d(x_{0},p_{i}))_{i=1,...,\ell}$, where $d$ computes distance in $\mathbb{R}^{2}$. The recurrent input to the RNN is a sequence of two-dimensional velocity vectors $\{v_{1},v_{2},...,v_{seq\textunderscore length}\}$, which assemble into a trajectory in the arena. The final hidden state of the RNN is mapped linearly back to $\mathbb{R}^{\ell}$, and softmax is applied. The loss term is the cross entropy of the output with $\operatorname{softmax}(d(x_{seq\textunderscore length},p_{i}))_{i=1,...,\ell}$, where $x_{seq\textunderscore length}=x_{0}+\sum_{i}v_{i}$ is the final position of the trajectory. We illustrate this set-up in Figure 2, and in greater detail in Supplemental Figure4. We then applied a Bayesian decoder to infer most likely position in the arena, based on the activations in $\mathbb{R}^{\ell}$, and report the distance between the predicted and true position. Untrained models have a mean error of $\approx 100$ cm in arenas of this size. Training data in the form of trajectories can be freely generated for the RNN, which alleviates the problem of overfitting. This allows us to study the effect of regularization on the sparsity of the underlying model, isolated from the other benefits of regularization. Hyperparameters for our experiments are listed in Appendix A.1.
Figure 2: The representation of the state transitions in the RNN. Left: ground truth of the navigation problem. Right: The internal representation used by a (toroidally regularized) RNN. Top: initialization in the ground truth places the agent at a random place in a box. The RNN uses a Gaussian tiling to represent the initial position. Bottom: In the ground truth, a sequence of velocity inputs moves the agent to a new position. In the RNN, this is represented by changing the hidden state to reflect the new position. In the figure, this is depicted by translation on a torus.
@@
Torus trained Torus fixed $\mathbb{R}^{3}$ trained $\mathbb{R}^{3}$ fixed
0% Sparse 4.79 (0.10) 4.67 (0.07) 5.85 (0.09) 4.77 (0.09)
75% Sparse 4.60 (0.06) 4.86 (0.30) 6.59 (0.85) 34.78 (42.10)
75% Lottery 8.20 (4.16) 7.28 (0.69) 35.44 (39.83) 60.37 (35.54)
90% Sparse 5.87 (0.55) 5.69 (0.43) 5.81 (0.45) 67.42 (33.27)
90% Lottery 15.72 (3.94) 12.49 (2.03) 20.67 (5.68) 46.56 (7.69)
@@
Sensitivity analysis on the regularizing factor ($\lambda$ in Equation (5)) can be found in Appendix G. This RNN architecture was originally introduced by [49] to study grid cells, certain specialized neurons in the entorhinal cortex which provide an internal map of space [15, 37, 55]. After the discovery of grid cells in 2005 [19], several authors [2, 18] show that toroidal neural connectivity can give rise to grid cell structures. However, to the best of our knowledge, no connectivity structures outside of the torus and $\mathbb{R}^{2}$ have been seriously investigated—the hand construction of these networks is time consuming, and small perturbations of weights can damage their effectiveness [53]. One surprising application of our moduli regularization is then the great ease with which it shows the learning advantages of the torus over arbitrary spaces, such as the sphere, for neural attractor dynamics on navigation. The success of toroidal, circular, and Klein bottle sparsification suggests that moduli regularization may provide a useful lens for exploratory work on highly structured neural connectivity. We expected these regularizers to outperform the sphere and 6-torus because $\mathbb{R}^{2}$ is a covering space (see summary in Appendix C) for the Klein bottle and the torus, making them apt compact representations for two-dimensional space. Since the torus is isomorphic to the product of two circles, a circular moduli space could also plausibly recreate toroidal connectivity.
4.2 Natural language processing
Natural language processing lacks clear a priori geometry to exploit [6]. Indeed, local stability of the hidden state is not a common hypothesis for NLP architectures like long-short term memory [22]. We test an RNN with three recurrent layers trained to predict subsequent words in a sequence on the Wikitext-2 dataset of high-quality Wikipedia articles [33]. Our implementation is based on the implementation of [3], which is a representative implementation of natural language RNNs. Hyperparameters for our experiments are listed in Appendix A.2.
Table 4: Mean test perplexity (standard deviation) of next-word prediction RNNs at 0%, 90%, and 98% sparsity, and testing the lottery ticket hypothesis with trained sparse architectures. $\,{}^{*}$ One outlier experiment with the $L_{1}$ regularizer, marked by a star, suffered representation collapse and had final perplexity $\approx 3.4\times 10^{10}$. We excluded this experiment from reported averages.
Torus No reg. $L_{1}$ Shuffled
0% Sparse 69.26 (16.31) 114.73 (1.56) 172.19 (94.41) 56.52 (5.64)
90% Sparse 43.41 (7.53) 91.12 (34.21) 118.18 (4.03) 67.98 (34.00)
90% Lottery 54.80 (8.70) 113.93 (0.92) 114.41 (0.92) 75.17 (21.55)
98% Sparse 55.80 (13.96) 118.45 (1.59) $98.78(31.97)^{*}$ 51.06 (10.15)
98% Lottery 61.44 (6.06) 115.21 (1.27) 115.06 (1.88) 62.58 (5.57)
Trained embeddings: As in navigation, we studied trained embeddings on the torus and in $\mathbb{R}^{3}$. To our surprise, training the embeddings resulted in very similar results to the navigation experiment, showing improvements in $\mathbb{R}^{3}$, but performing poorly on the torus. We give results of our experiments in Table 5. We suspect this deficiency might reflect the regularizer abetting overfitting in the trained case, though we did not extensively investigate this hypothesis.
Table 5: Mean perplexity (standard deviation) of five experiments at 0%, 90%, and 98% sparsity in a next-word-predictor RNN. Training embeddings as we did in this context can be actively harmful to the final model quality.
@@ 4.2 Natural language processing
0% Sparse 119.90 (5.92) 69.26 (16.31) 109.49 (19.83) 116.18 (1.97)
90% Sparse 45.70 (7.29) 40.35 (4.30) 68.53 (19.21) 117.25 (1.54)
90% Lottery 109.56 (5.96) 56.43 (2.71) 114.04 (0.37) 114.56 (0.00)
98% Sparse 107.24 (24.92) 65.53 (9.40) 62.52 (10.46) 121.37 (4.02)
98% Lottery 115.36 (0.65) 62.57 (7.54) 107.92 (10.52) 116.37 (0.67)
@@
Table 6: Mean test perplexity (standard deviation) of next-word prediction RNNs at baseline, 90%, and 98% sparsity, trained with a variety of moduli regularizers, and testing the lottery ticket hypothesis. Underlined values are ab initio sparse models that outperform their sparsified antecedents.
@@
0% Sparse 69.26 (16.31) 58.03 (15.78) 65.51 (6.48) 64.35 (10.89) 57.17 (12.80)
90% Sparse 43.41 (7.53) 49.33 (10.61) 49.38 (10.76) 51.03 (10.46) 44.91 (3.63)
90% Lottery 54.80 (8.70) 51.32 (11.08) 55.52 (11.56) 57.78 (10.87) 46.16 (6.70)
98% Sparse 55.80 (13.96) 44.88 (3.32) 52.86 (12.57) 58.86 (7.44) 57.68 (10.61)
98% Lottery 61.44 (6.06) 63.84 (3.69) 57.89 (5.79) 53.34 (2.18) 63.73 (5.78)
Table 7: Mean test perplexity (standard deviation) on RNNs with trained embeddings. Trained embeddings in the torus allows the model to overfit, while it improves the moduli approximation in $\mathbb{R}^{3}$.
R3 trained R3 fixed Torus trained Torus fixed
90% Sparse 81.50 (27.70) 106.38 (21.23) 103.22 (27.56) 43.41 (7.53)
90% Lottery 101.93 (15.26) 114.25 (0.39) 108.59 (12.42) 54.80 (8.70)
@@
Our results show that moduli regularization remains a highly effective sparsification method in non-geometric RNNs. Moduli regularizers (Table 6) improve the accuracy of trained NLP models compared to both non-regularized and $L_{1}$ regularized models (Table 4). While they do not consistently outperform random sparsification at 98% sparsity, some moduli regularizers create sparse architectures that are more resilient to retraining with randomized weights, in some cases defying the lottery ticket hypothesis [9].
4.3 The adding problem
The adding problem is an experiment to study long term memory in recurrent architectures (see [27], Section 4.1, or [28], Section 5.1). The input data is a sequence of numbers chosen uniformly at random between 0 and 1, as well as a mask which specifies 1 at two points in the sequence, and 0 elsewhere. The target for the RNN is the sum of the product of the input data row with the mask row. We illustrate an example of the adding problem in Figure 3. Hyperparameters for our experiments are listed in Appendix A.3, and all inhibitor functions are the difference of Gaussians (Appendix E) unless otherwise stated. In the adding problem, always predicting the value 1 has mean square error .$\approx.17$, and root mean square error (RMSE) .$\approx.41$; a model has learned if it substantially improves on these baselines.
input .2 .6 .9 .1 .4 .4 .7 .1 .8 output
mask 0 0 1 0 1 0 0 0 0 1.3
Figure 3: The adding problem inputs (left) and intended output (right).
We tested a single layer RNN with a variety of moduli regularizers on this task to experiment on the manner in which moduli regularizers interact with long term memories. Moduli regularization broadly generalizes previous specialized architectures designed for the adding problem: for example, IndRNNs [28] are equivalent to RNNs trained with moduli regularization, where the inhibitor function $f$ is given by $f(0)=0,$ and $f(x)=\infty$ for all $x>0$.
Control experiments: We repeat the control experiments used in the Navigation and control experiments. Table 8 shows that toroidal moduli regularizers outperform both $L_{1}$ regularizers and shuffled moduli regularizers at the 95% sparsity threshold.
Table 8: Root mean square error (standard deviation) over five experiments at 95% sparsity. Toroidal moduli regularization results in the highest quality sparse model, though no lottery ticket sparse architectures are able to train.
Torus No reg L1 Permuted
0% Sparse RMSE 0.040 (0.006) 0.036 (0.006) 0.036 (0.007) 0.042 (0.010)
95% Sparse RMSE 0.111 (0.035) 0.142 (0.122) 0.265 (0.113) 0.142 (0.038)
95% Lottery RMSE 0.404 (0.003) 0.407 (0.001) 0.409 (0.004) 0.400 (0.010)
Trained embeddings: Training the embedding allows the model to find superior regularizer arrangements for both the torus and $\mathbb{R}^{3}$, as we show in Table 9.
Torus trained Torus fixed $\mathbb{R}^{3}$ trained $\mathbb{R}^{3}$ fixed
0% Sparse RMSE 0.035 (0.005) 0.040 (0.006) 0.044 (0.005) 0.037 (0.005)
95% Sparse RMSE 0.055 (0.012) 0.111 (0.035) 0.044 (0.012) 0.050 (0.026)
95% Lottery RMSE 0.406 (0.004) 0.404 (0.003) 0.407 (0.002) 0.406 (0.002)
Table 9: Root mean square error (standard deviation) over five experiments at 95% sparsity. Training the embedding results in improvements in both toroidal and $\mathbb{R}^{3}$ regimes.
Moduli spaces: We compare the same moduli space hyperparameters as tested above in Table 10. The torus, circle, and 6-torus perform well at 95% sparsity, while the Klein bottle and sphere perform poorly.
Table 10: Root mean square error (standard deviation) over five experiments at 95% sparsity, with lottery ticket trained sparse architectures.
Torus Circle Sphere Klein 6-Torus
0% Sparse RMSE 0.040 (0.006) 0.038 (0.005) 0.042 (0.014) 0.039 (0.007) 0.037 (0.009)
95% Sparse RMSE 0.111 (0.035) 0.097 (0.016) 0.148 (0.114) 0.214 (0.122) 0.088 (0.006)
95% Lottery RMSE 0.404 (0.003) 0.407 (0.002) 0.404 (0.005) 0.404 (0.004) 0.353 (0.088)
Our results show that moduli regularization is compatible with long term recall in RNNs, an important feature for implementation. Moreover, it is important as a sample use case where learned moduli spaces substantially outperform fixed ones.